# Inference for Deep Neural Network Estimators in Generalized Nonparametric Models

### v3

@@ 1 Introduction
Modern inference approaches include distribution-free conformal estimation inference (Lei et al., 2018; Angelopoulos & Bates, 2021; Huang et al., 2024), extensions of jackknife inference within conformal estimation (Alaa & Van Der Schaar, 2020; Kim et al., 2020; Barber et al., 2021), debiasing methods (Athey et al., 2018; Guo et al., 2021), residual-based bootstrap techniques (Zhang & Politis, 2023), and the use of $U$ statistics in regressions, random forests and neural networks (Mentch & Hooker, 2016; Wager & Athey, 2018; Schupbach et al., 2020; Wang & Wei, 2022; Fei & Li, 2024). Several distinctions and limitations remain. For example, conformal prediction constructs distribution-free prediction sets for $y$, targeting outcome coverage rather than inference on $\mathbb{E}(y|\mathbf{x})$. Its intervals thus absorb both outcome noise and model uncertainty, making them conservative and inefficient for mean inference. Clinically, inference on risk classifications derived from patients’ risk profiles, which are based on the estimated conditional mean (i.e., risk score), is more actionable, as medical decisions (e.g., triage or treatment intensity) rely on risk stratifications rather than prediction sets (Lundberg et al., 2018; Van Calster et al., 2019). Jackknife and related leave-one-out methods assume stable, nearly independent residuals, conditions violated by DNNs whose estimates depend heavily on individual samples. Debiasing approaches (Athey et al., 2018; Guo et al., 2021) and residual bootstraps (Zhang & Politis, 2023) remain limited to linear models, while resampling-based inference for high-dimensional GLMs (Fei & Li, 2021) is fully parametric. Recent work on uncertainty quantification for random forests (Mentch & Hooker, 2016; Wager & Athey, 2018) and neural networks (Schupbach et al., 2020; Fei & Li, 2024) offers valuable insights but remains largely confined to continuous outcomes, leaving inference for GNRMs such as logistic and Poisson regressions unresolved.
@@
Theoretically, for valid inference on $\mathbb{E}(y|\mathbf{x})$, we need to establish the convergence rate of neural networks in a GNRM framework. Since the universal approximation theorem (Hornik et al., 1989), there has been much progress in understanding the convergence rates of neural networks (McCaffrey & Gallant, 1994; Kohler & Krzyżak, 2005; Yarotsky, 2017; Shen et al., 2019; Kohler & Langer, 2021; Lu et al., 2021; Shen et al., 2021; Jiao et al., 2023; Fan et al., 2024; Wang et al., 2024; Bhattacharya et al., 2024; Yan et al., 2025). For instance, Shen et al. (2021) established the convergence rate under nonparametric regression assuming the $p$-th moment of response is bounded for some $p>1$; Fan et al. (2024) explored nonasymptotic error bounds of the estimators under several loss functions, such as the least-absolute deviation loss and Huber loss; Wang et al. (2024) proposed an efficient and robust nonparametric regression estimator based on an estimated maximum likelihood estimator; Fan & Gu (2024); Bhattacharya et al. (2024) investigated deep neural network estimators for nonparametric interaction models with diverging input dimensions, focusing on theoretical properties such as minimax optimality, approximation error control, and sparsity-driven regularization. However, quantifying uncertainty of DNN-estimated $\mathbb{E}(y|\mathbf{x})$ within the framework of GNRM introduces additional challenges. The existing approaches typically assume a nonparametric regression framework, where the distribution of $y-\mathbb{E}(y|\mathbf{x})$ does not depend on $\mathbf{x}$. This assumption simplifies the theoretical analysis and the derivation of convergence results (Takezawa, 2005; Györfi et al., 2006; Fan, 2018). In generalized nonparametric regression settings, however, the assumption often fails due to heteroskedasticity, as the distribution of $y-\mathbb{E}(y|\mathbf{x})$ may still depend on $\mathbf{x}$. Consequently, the standard techniques for bounding error terms and controlling variance no longer apply, and the established results on uniform convergence or concentration inequalities (Bartlett et al., 2020) are invalid. As a result, it is unclear whether these methods are applicable for drawing inference on GNRMs estimated by DNNs.
@@
Our methodological and theoretical contributions are as follows. First, we propose a DNN estimator for subject specific means under GNRMs and develop an inference framework that explicitly accounts for covariate-dependent error distributions induced by heteroskedasticity. We provide new proof techniques, including new criterion loss function, covering-number based concentration and a truncation strategy that handles exponential-tailed, heteroskedastic noise, improving the results of Schmidt-Hieber (2020). For example, with the general loss induced by GNRMs, the quadratic-based arguments of Schmidt-Hieber (2020) no longer apply. This necessitates a complete redevelopment of the error bounds, derived instead from the intrinsic curvature and Lipschitz properties in the new GNRM setting. As a result, all estimation error analyses are novel and move beyond Gaussian approximations. Additionally, the presence of exponential tailed, heteroskedastic noise complicates uniform convergence, which we address through a truncation strategy. The truncation scheme partitions the analysis into bounded and unbounded noise regimes, where standard convergence arguments apply in the former and concentration inequalities control the latter, yielding sharper error bounds. These theoretical tools enable us to extend the continuous outcome setting of Schmidt-Hieber (2020), accommodating both continuous and discrete outcomes, and thereby generalizing their framework to the broader GNRM class. Second, to quantify the uncertainty of the estimates, we propose an ensemble subsampling method (ESM) under GNRMs, developed from U-statistics (Frees, 1989; Hoeffding, 1992; Lee, 2019; Boroskikh, 2020) and Hoeffding decomposition (Hoeffding, 1992). Our model-free variance estimator leverages the intrinsic properties of Hoeffding decomposition and accommodates varying population levels as well as individual variance structures. Moreover, we employ incomplete U-statistics to reduce computational cost, enabling the resample size to grow proportionally with the sample size while limiting the total number of resamples. To mitigate the resulting Monte Carlo bias, we refine the variance estimator of Wager & Athey (2018) for bias correction.
@@ 2.1 Deep Neural Networks
with the recursive expression $f_{l}(\mathbf{x})=\sigma(\mathbf{W}_{l-1}f_{l-1}(\mathbf{x})+\mathbf{v}_{l-1}),\quad f_{1}(\mathbf{x})=\sigma(\mathbf{W}_{0}\mathbf{x}+\mathbf{v}_{0}).$ Here, $\mathbf{W}_{l}\in\mathbb{R}^{p_{l+1}\times p_{l}}$ and vectors $\mathbf{v}_{l}\in\mathbb{R}^{p_{l+1}}$ ($l\in[L]$) are the parameters of this DNN, and $\sigma:\mathbb{R}\to\mathbb{R}$ is the activation function, applied element-wise to vectors. A commonly used activation function is the rectified linear unit (ReLU) function (Nair & Hinton, 2010), or $\sigma(x)=\max(x,0).$ We focus on ReLU due to its piecewise linearity and projection properties, with potential extensions to other activation functions as discussed in Section 6.
@@ 2.2 Generalized Nonparametric Regression
Consider outcomes from the exponential family, including commonly used distributions such as normal, Bernoulli, Poisson, and binomial. The density (or mass) function for a random variable $y$ in the (single-parameter) exponential family is $p(y|\theta)=h(y)\exp\left\{\theta y-\psi(\theta)\right\},$ where $\theta$ is the canonical parameter, $0<h(y)<\infty$ ensures normalization, and $\psi(\theta)$ is convex and smooth, satisfying $\mathbb{E}(y|\theta)=\psi^{\prime}(\theta)$, where $\psi^{\prime}(\cdot)$ denotes the derivative of $\psi(\cdot)$ and the inverse function of $\psi^{\prime}(\cdot)$ is known as the link function (Brown, 1986; Fei & Li, 2021).
@@
Nonparametric Binomial Regression: Given covariates $\mathbf{x}$, the response variable $y$ follows a Binomial distribution with a fixed number of trials $n_{\text{trial}}$ and the success probability $p(\mathbf{x})=\frac{1}{1+e^{-f_{0}(\mathbf{x})}}$. This corresponds to $h(y)=\binom{n_{\text{trial}}}{y}$ and $\psi(\theta)=n_{\text{trial}}\log(1+e^{\theta})$. In this setting, the estimator (2.4) is specified as
@@ 2.3 Empirical Risk Minimization and Function Smoothness
where we use $\mathbb{E}[y|\mathbf{x}]=\psi^{\prime}(f_{0}(\mathbf{x}))$. This equals the Bregman divergence between $f(\mathbf{x})$ and $f_{0}(\mathbf{x})$, which is nonnegative and vanishes if and only if $f(\mathbf{x})=f_{0}(\mathbf{x})$. We then quantify the estimation error of $\widehat{f}$, any estimator of $f_{0}$, by evaluating the pointwise estimation error at an independent test point $\bm{X}\sim\mathbf{x}_{1}$: $\ell(\bm{X};\widehat{f},f_{0})=-\psi^{\prime}(f_{0}(\bm{X}))(\widehat{f}(\bm{X})-f_{0}(\bm{X}))+\psi(\widehat{f}(\bm{X}))-\psi(f_{0}(\bm{X})).$ We further define the average estimation error, our main theoretical target, as
@@
Computing the exact minimizer $\widehat{f}_{n}^{{\mathop{\mathrm{opt}}}}$ in (2.4) is challenging due to the nonconvex loss (Fan & Gu, 2024). Instead, we obtain an approximate minimizer $\widehat{f}_{n}$ using optimization algorithms such as gradient descent. To quantify the optimization error, we introduce
@@
For desirable properties of $\widehat{f}_{n}$, we assume $f_{0}$ belongs to a Hölder class (Schmidt-Hieber, 2020; Fan & Gu, 2024) with parameters $\gamma$, $K>0$, and domain $\mathbb{D}\subset\mathbb{R}^{r}$:
@@
where $\lfloor\gamma\rfloor$ is the largest integer strictly smaller than $\gamma$, and $\partial^{{\bm{\beta}}}=\partial^{\beta_{1}}\ldots\partial^{\beta_{r}}$ with ${\bm{\beta}}=(\beta_{1},...,\beta_{r})^{\top}\in\mathbb{N}^{r}$. We further assume that $f_{0}$ is a $(q+1)$-composition ($q\in\mathbb{N}$) of several Hölder functions. That is, for some vectors $\mathbf{d}=(d_{0},...,d_{q+1})\in\mathbb{N}_{+}^{q+2}$, $\mathbf{t}=(t_{0},...,t_{q})\in\mathbb{N}_{+}^{q+1}$ and ${\bm{\gamma}}=(\gamma_{0},...,\gamma_{q})\in\mathbb{R}_{+}^{q+1}$, $f_{0}\in\mathcal{G}(q,\mathbf{d},\mathbf{t},{\bm{\gamma}},K)$, where
$\mathcal{G}(q,\mathbf{d},\mathbf{t},{\bm{\gamma}},K):=\bigg\{$ $f=g_{q}\circ\cdots\circ g_{0}:g_{i}=(g_{ij})_{j}:[a_{i},b_{i}]^{d_{i}}\to[a_{i+1},b_{i+1}]^{d_{i+1}},$
$g_{ij}\in\mathcal{G}_{t_{i}}^{\gamma_{i}}([a_{i},b_{i}]^{t_{i}})\text{ with }|a_{i}|,|b_{i}|\leq K\bigg\}.$
The motivation for the assumptions on $f_{0}$ is that the Hölder class provides a flexible framework for quantifying smoothness, capturing both integer and fractional orders, and serves as a natural benchmark for assessing neural network approximation (Yarotsky, 2017), while the compositional structure mirrors the architectural design of deep neural networks with each layer applying a transformation to the output of the previous one (Yarotsky, 2017). We also note that $\mathbf{d}$ and $\mathbf{t}$ characterize the dimensions of the function, while $\bm{\gamma}$} represents the measure of smoothness. For instance, if
@@
and $g_{ij}$ are twice continuously differentiable, then smoothness ${\bm{\gamma}}=(2,2,2)$, dimensions $\mathbf{d}=(24,6,3,1)^{\top}$ and $\mathbf{t}=(4,2,3)^{\top}$. We further define $\gamma_{j}^{*}=\gamma_{j}\Pi_{l=i+1}^{q}(\gamma_{l}\wedge 1)$ and $\phi_{n}=\max_{j=0,...,q}n^{-2\gamma_{j}^{*}/(2\gamma_{j}^{*}+t_{j})}$.
@@ 2.4 Ensemble Subsampling Method (ESM)
Consider a scenario where $\mathbf{x}$ is drawn from an unknown probability measure $\mathbb{P}_{\mathbf{x}}$ supported on $\mathcal{X}$, and for an input $\mathbf{x}_{{\mathop{\mathrm{new}}}}=\mathbf{x}_{*}\in\mathcal{X}$, the unobserved outcome $y_{{\mathop{\mathrm{new}}}}$ follows (2.3). Our goal is to estimate $\mathbb{E}(y_{{\mathop{\mathrm{new}}}}|\mathbf{x}_{{\mathop{\mathrm{new}}}}=\mathbf{x}_{*})=\psi^{\prime}(f_{0}(\mathbf{x}_{*}))$ and its uncertainty. For this, we introduce the Ensemble Subsampling Method (ESM), based on subsampling techniques (Wager & Athey, 2018; Fei & Li, 2024), to construct ensemble estimators and confidence intervals. ESM consists of the following components, which are summarized in Figure 1 as well.
@@
    where $\Delta^{{\mathop{\mathrm{opt}}}}_{b}$ quantifies the optimization error (Fan & Gu, 2024), and $\widehat{f}^{b,{\mathop{\mathrm{opt}}}}$ and $\Delta_{r}(\widehat{f}^{b},\widehat{f}^{b,{\mathop{\mathrm{opt}}}})$ are defined respectively in (2.4) and (2.6) on the subsample indexed by $\mathcal{I}^{b}$.
@@
The distribution of $\widehat{f}^{b}(\mathbf{x}_{*})$ for a fixed $\mathbf{x}_{*}$ is intractable. However, since $\widehat{f}^{b}(\mathbf{x}_{*})$ is permutation symmetric with respect to $\mathcal{I}^{b}$, the ensemble estimator $\widehat{f}^{B^{*}}(\mathbf{x}_{*})$ forms a generalized U-statistic, which we later show to be asymptotically normal by leveraging U-statistics theory. In practice, we allow $r$ to grow with $n$, but evaluating all $B^{*}={\binom{n}{r}}$ models becomes computationally infeasible. To overcome this, we use a stochastic approximation by randomly drawing $B$ size-$r$ subsamples from $\mathcal{D}_{n}$. This involves independently sampling indices $b_{1},\dots,b_{B}$ from $1$ to $\binom{n}{r}$}, where each subsample is indexed by $\mathcal{I}^{b_{j}}$, and computing

### v4

@@ 1 Introduction
Modern inference approaches include distribution-free conformal estimation inference (Lei et al., 2018; Angelopoulos and Bates, 2021; Huang et al., 2024), extensions of jackknife inference within conformal estimation (Alaa and Van Der Schaar, 2020; Kim et al., 2020; Barber et al., 2021), debiasing methods (Athey et al., 2018; Guo et al., 2021), residual-based bootstrap techniques (Zhang and Politis, 2023), and the use of $U$ statistics in regressions, random forests and neural networks (Mentch and Hooker, 2016; Wager and Athey, 2018; Schupbach et al., 2020; Wang and Wei, 2022; Fei and Li, 2024). Several distinctions and limitations remain. For example, conformal prediction constructs distribution-free prediction sets for $y$, targeting outcome coverage rather than inference on $\mathbb{E}(y|\mathbf{x})$. Its intervals thus absorb both outcome noise and model uncertainty, making them conservative and inefficient for mean inference. Clinically, inference on risk classifications derived from patients’ risk profiles, which are based on the estimated conditional mean (i.e., risk score), is more actionable, as medical decisions (e.g., triage or treatment intensity) rely on risk stratifications rather than prediction sets (Lundberg et al., 2018; Van Calster et al., 2019). Jackknife and related leave-one-out methods assume stable, nearly independent residuals, conditions violated by DNNs whose estimates depend heavily on individual samples. Debiasing approaches (Athey et al., 2018; Guo et al., 2021) and residual bootstraps (Zhang and Politis, 2023) remain limited to linear models, while resampling-based inference for high-dimensional GLMs (Fei and Li, 2021) is fully parametric. Recent work on uncertainty quantification for random forests (Mentch and Hooker, 2016; Wager and Athey, 2018) and neural networks (Schupbach et al., 2020; Fei and Li, 2024) offers valuable insights but remains largely confined to continuous outcomes, leaving inference for GNRMs such as logistic and Poisson regressions unresolved.
@@
Theoretically, for valid inference on $\mathbb{E}(y|\mathbf{x})$, we need to establish the convergence rate of neural networks in a GNRM framework. Since the universal approximation theorem (Hornik et al., 1989), there has been much progress in understanding the convergence rates of neural networks (McCaffrey and Gallant, 1994; Kohler and Krzyżak, 2005; Yarotsky, 2017; Shen et al., 2019; Kohler and Langer, 2021; Lu et al., 2021; Shen et al., 2021; Jiao et al., 2023; Fan et al., 2024; Wang et al., 2024; Bhattacharya et al., 2024; Yan et al., 2025). For instance, Shen et al. (2021) established the convergence rate under nonparametric regression assuming the $p$-th moment of response is bounded for some $p>1$; Fan et al. (2024) explored nonasymptotic error bounds of the estimators under several loss functions, such as the least-absolute deviation loss and Huber loss; Wang et al. (2024) proposed an efficient and robust nonparametric regression estimator based on an estimated maximum likelihood estimator; Fan and Gu (2024); Bhattacharya et al. (2024) investigated deep neural network estimators for nonparametric interaction models with diverging input dimensions, focusing on theoretical properties such as minimax optimality, approximation error control, and sparsity-driven regularization. However, quantifying uncertainty of DNN-estimated $\mathbb{E}(y|\mathbf{x})$ within the framework of GNRM introduces additional challenges. The existing approaches typically assume a nonparametric regression framework, where the distribution of $y-\mathbb{E}(y|\mathbf{x})$ does not depend on $\mathbf{x}$. This assumption simplifies the theoretical analysis and the derivation of convergence results (Takezawa, 2005; Györfi et al., 2006; Fan, 2018). In generalized nonparametric regression settings, however, the assumption often fails due to heteroskedasticity, as the distribution of $y-\mathbb{E}(y|\mathbf{x})$ may still depend on $\mathbf{x}$. Consequently, the standard techniques for bounding error terms and controlling variance no longer apply, and the established results on uniform convergence or concentration inequalities (Bartlett et al., 2020) are invalid. As a result, it is unclear whether these methods are applicable for drawing inference on GNRMs estimated by DNNs.
@@
Our methodological and theoretical contributions are as follows. First, we propose a DNN estimator for subject specific means under GNRMs and develop an inference framework that explicitly accounts for covariate-dependent error distributions induced by heteroskedasticity. We provide new proof techniques, including new criterion loss function, covering-number based concentration and a truncation strategy that handles exponential-tailed, heteroskedastic noise in Section D in the Supplementary Material, improving the results of Schmidt-Hieber (2020). For example, with the general loss induced by GNRMs, the quadratic-based arguments of Schmidt-Hieber (2020) no longer apply. This necessitates a complete redevelopment of the error bounds, derived instead from the intrinsic curvature and Lipschitz properties in the new GNRM setting. As a result, all estimation error analyses are novel and move beyond Gaussian approximations. Additionally, the presence of exponential tailed, heteroskedastic noise complicates uniform convergence, which we address through a truncation strategy. The truncation scheme partitions the analysis into bounded and unbounded noise regimes, where standard convergence arguments apply in the former and concentration inequalities control the latter, yielding sharper error bounds. These theoretical tools enable us to extend the continuous outcome setting of Schmidt-Hieber (2020), accommodating both continuous and discrete outcomes, and thereby generalizing their framework to the broader GNRM class. Second, to quantify the uncertainty of the estimates, we propose an ensemble subsampling method (ESM) under GNRMs, developed from U-statistics (Frees, 1989; Hoeffding, 1992; Lee, 2019; Boroskikh, 2020) and Hoeffding decomposition (Hoeffding, 1992). Our model-free variance estimator leverages the intrinsic properties of Hoeffding decomposition and accommodates varying population levels as well as individual variance structures. Moreover, we employ incomplete U-statistics to reduce computational cost, enabling the resample size to grow proportionally with the sample size while limiting the total number of resamples. To mitigate the resulting Monte Carlo bias, we refine the variance estimator of Wager and Athey (2018) for bias correction.
@@ 2.1 Deep Neural Networks
with the recursive expression $f_{l}(\mathbf{x})=\sigma(\mathbf{W}_{l-1}f_{l-1}(\mathbf{x})+\mathbf{v}_{l-1}),\quad f_{1}(\mathbf{x})=\sigma(\mathbf{W}_{0}\mathbf{x}+\mathbf{v}_{0}).$ Here, $\mathbf{W}_{l}\in\mathbb{R}^{p_{l+1}\times p_{l}}$ and vectors $\mathbf{v}_{l}\in\mathbb{R}^{p_{l+1}}$ ($l\in[L]$) are the parameters of this DNN, and $\sigma:\mathbb{R}\to\mathbb{R}$ is the activation function, applied element-wise to vectors. A commonly used activation function is the rectified linear unit (ReLU) function (Nair and Hinton, 2010), or $\sigma(x)=\max(x,0).$ We focus on ReLU due to its piecewise linearity and projection properties, with potential extensions to other activation functions as discussed in Section 6.
@@ 2.2 Generalized Nonparametric Regression
Consider outcomes from the exponential family, including commonly used distributions such as normal, Bernoulli, Poisson, and binomial. The density (or mass) function for a random variable $y$ in the (single-parameter) exponential family is $p(y|\theta)=h(y)\exp\left\{\theta y-\psi(\theta)\right\},$ where $\theta$ is the canonical parameter, $0<h(y)<\infty$ ensures normalization, and $\psi(\theta)$ is convex and smooth, satisfying $\mathbb{E}(y|\theta)=\psi^{\prime}(\theta)$, where $\psi^{\prime}(\cdot)$ denotes the derivative of $\psi(\cdot)$ and the inverse function of $\psi^{\prime}(\cdot)$ is known as the link function (Brown, 1986; Fei and Li, 2021).
@@
Nonparametric Binomial Regression: Given covariates $\mathbf{x}$, the response variable $y$ follows a Binomial distribution with a fixed number of trials $n_{\text{trial}}$ and the success probability $p(\mathbf{x})=\frac{1}{1+e^{-f_{0}(\mathbf{x})}}$. This corresponds to $h(y)={n_{\text{trial}}\choose{y}}$ and $\psi(\theta)=n_{\text{trial}}\log(1+e^{\theta})$. In this setting, the estimator (2.4) is specified as
@@ 2.3 Empirical Risk Minimization and Function Smoothness
where we use $\mathbb{E}[y|\mathbf{x}]=\psi^{\prime}(f_{0}(\mathbf{x}))$. This equals the Bregman divergence between $f(\mathbf{x})$ and $f_{0}(\mathbf{x})$, which is nonnegative and vanishes if and only if $f(\mathbf{x})=f_{0}(\mathbf{x})$. We then quantify the estimation error of $\widehat{f}$, any estimator of $f_{0}$, by evaluating the pointwise estimation error at an independent test point $\bm{X}$: $\ell(\bm{X};\widehat{f},f_{0})=-\psi^{\prime}(f_{0}(\bm{X}))(\widehat{f}(\bm{X})-f_{0}(\bm{X}))+\psi(\widehat{f}(\bm{X}))-\psi(f_{0}(\bm{X})).$ We further define the average estimation error, our main theoretical target, as
@@
Computing the exact minimizer $\widehat{f}_{n}^{{\mathop{\mathrm{opt}}}}$ in (2.4) is challenging due to the nonconvex loss (Fan and Gu, 2024). Instead, we obtain an approximate minimizer $\widehat{f}_{n}$ using optimization algorithms such as gradient descent. To quantify the optimization error, we introduce
@@
For desirable properties of $\widehat{f}_{n}$, we assume $f_{0}$ belongs to a Hölder class (Schmidt-Hieber, 2020; Fan and Gu, 2024) with parameters $\gamma$, $K>0$, and domain $\mathbb{D}\subset\mathbb{R}^{r}$:
@@
where $\lfloor\gamma\rfloor$ is the largest integer strictly smaller than $\gamma$, and $\partial^{{\bm{\beta}}}=\partial^{\beta_{1}}\ldots\partial^{\beta_{r}}$ with ${\bm{\beta}}=(\beta_{1},...,\beta_{r})^{\top}\in\mathbb{N}^{r}$. We further assume that $f_{0}$ is a $(q+1)$-composition ($q\in\mathbb{N}$) of several Hölder functions. That is, for some vectors $\mathbf{d}=(d_{0},...,d_{q+1})\in\mathbb{N}_{+}^{q+2}$, $\mathbf{t}=(t_{0},...,t_{q})\in\mathbb{N}_{+}^{q+1}$ and $\mathbf{m}=(m_{0},...,m_{q})\in\mathbb{R}_{+}^{q+1}$, $f_{0}\in\mathcal{G}(q,\mathbf{d},\mathbf{t},\mathbf{m},K)$, where
$\mathcal{G}(q,\mathbf{d},\mathbf{t},\mathbf{m},K):=\bigg\{$ $f=g_{q}\circ\cdots\circ g_{0}:g_{i}=(g_{ij})_{j}:[a_{i},b_{i}]^{d_{i}}\to[a_{i+1},b_{i+1}]^{d_{i+1}},$
$g_{ij}\in\mathcal{G}_{t_{i}}^{m_{i}}([a_{i},b_{i}]^{t_{i}})\text{ with }|a_{i}|,|b_{i}|\leq K\bigg\}.$
The motivation for the assumptions on $f_{0}$ is that the Hölder class provides a flexible framework for quantifying smoothness, capturing both integer and fractional orders, and serves as a natural benchmark for assessing neural network approximation (Yarotsky, 2017), while the compositional structure mirrors the architectural design of deep neural networks with each layer applying a transformation to the output of the previous one (Yarotsky, 2017). We also note that $\mathbf{d}$ and $\mathbf{t}$ characterize the dimensions of the function, while $\mathbf{m}$ represents the measure of smoothness. For instance, if
@@
and $g_{ij}$ are twice continuously differentiable, then smoothness $\mathbf{m}=(2,2,2)$, dimensions $\mathbf{d}=(24,6,3,1)^{\top}$ and $\mathbf{t}=(4,2,3)^{\top}$. We further define $m_{j}^{*}=m_{j}\Pi_{l=i+1}^{q}(m_{l}\wedge 1)$ and $\phi_{n}=\max_{j=0,...,q}n^{-2m_{j}^{*}/(2m_{j}^{*}+t_{j})}$.
@@ 2.4 Ensemble Subsampling Method (ESM)
Consider a scenario where $\mathbf{x}$ is drawn from an unknown probability measure $\mathbb{P}_{\mathbf{x}}$ supported on $\mathcal{X}$, and for an input $\mathbf{x}_{{\mathop{\mathrm{new}}}}=\mathbf{x}_{*}\in\mathcal{X}$, the unobserved outcome $y_{{\mathop{\mathrm{new}}}}$ follows (2.3). Our goal is to estimate $\mathbb{E}(y_{{\mathop{\mathrm{new}}}}|\mathbf{x}_{{\mathop{\mathrm{new}}}}=\mathbf{x}_{*})=\psi^{\prime}(f_{0}(\mathbf{x}_{*}))$ and its uncertainty. For this, we introduce the Ensemble Subsampling Method (ESM), based on subsampling techniques (Wager and Athey, 2018; Fei and Li, 2024), to construct ensemble estimators and confidence intervals. ESM consists of the following components, which are summarized in Figure 1 as well.
@@
    where $\Delta^{{\mathop{\mathrm{opt}}}}_{b}$ quantifies the optimization error (Fan and Gu, 2024), and $\widehat{f}^{b,{\mathop{\mathrm{opt}}}}$ and $\Delta_{r}(\widehat{f}^{b},\widehat{f}^{b,{\mathop{\mathrm{opt}}}})$ are defined respectively in (2.4) and (2.6) on the subsample indexed by $\mathcal{I}^{b}$.
@@
The distribution of $\widehat{f}^{b}(\mathbf{x}_{*})$ for a fixed $\mathbf{x}_{*}$ is intractable. However, since $\widehat{f}^{b}(\mathbf{x}_{*})$ is permutation symmetric with respect to $\mathcal{I}^{b}$, the ensemble estimator $\widehat{f}^{B^{*}}(\mathbf{x}_{*})$ forms a generalized U-statistic, which we later show to be asymptotically normal by leveraging U-statistics theory. In practice, we allow $r$ to grow with $n$, but evaluating all $B^{*}={{n}\choose{r}}$ models becomes computationally infeasible. To overcome this, we use a stochastic approximation by randomly drawing $B$ size-$r$ subsamples from $\mathcal{D}_{n}$. This involves independently sampling indices $b_{1},\dots,b_{B}$ from $1$ to ${n}\choose{r}$}, where each subsample is indexed by $\mathcal{I}^{b_{j}}$, and computing